### v1

@@ Abstract
The key components of the framework include the utilization of baseline XAI methods (K methods), a ’Weighted Average’ explanation, a non-linear network (’Non-linear structure’), and an ’Up-Sampling’ unit. The framework incorporates a range of explanations obtained from established XAI methods and computes an adjusted baseline explanation (’Weighted Average’). The local explanations and the ’Weighted Average’ are concatenated and fed into a suitably designed non-linear structure (Swin-UNETR (a) and U-Net (b)). The reconstructed optimal explanation feeds an upsampling deconvolution layer to increase the resolution (high-resolution (HR) local explanation). The cost function evaluates the similarity between the weighted average and the low-resolution prediction, which is added to the faithfulness and complexity scores of the explanation. a)The 3D application uses images from the Top-Oslo dataset ([25]). b) The 2D application uses images from the CIFAR-10 dataset ([28]).
To encapsulate, in this study, we propose a framework aimed at providing optimal explanations for pre-trained deep networks in specific computer vision tasks. Characterized by high faithfulness and low complexity, this framework ensures accuracy and comprehension in the provided explanations. Additionally, we develop a separate framework designed to deliver explanations in high-resolution spaces, enhancing our ability to capture and understand significant feature contributions effectively. Our framework is tested in applications across diverse domains, including automobiles, animals, and neuroscience, encompassing both 2D and 3D multi-classification and binary classification tasks. Finally, we formalize a generalized framework capable of optimizing explanations for N-dimensional inputs, extending the utility of our approach beyond specific computer vision tasks to offer a versatile solution for interpreting complex data across various dimensions and domains. Our results suggest that optimal explanations based on specific criteria are indeed derivable and address the issue of inter-method variability in the current XAI literature.
Results
Classification training and optimization of explanations
The deep neural networks were trained for binary and multi-label classification tasks, in two and three dimensional spaces. Hyperparameter tuning was conducted using a range of learning rates () to optimize performance for both tasks. The sensitivity of the results for these different learning rates is illustrated with light blue shading in Extended Data Fig. 1 a,b and c,d. Results show the minimum, median and maximum loss and accuracy values that the deep networks achieved in the training and validation datasets for the 3D and 2D applications, respectively. In the 3D binary classification task, the network achieved approximately 75.0% validation accuracy and around 70.0% training accuracy (Extended Data Fig. 1a,b). For the 2D multi-classification task, the highest validation accuracy exceeded 80.0%, with training accuracy surpassing 75.0%. Extended Data Fig. 1a-d. demonstrate the effective training of Resnet-18 and ResNet-50 models on Top-OSLO and CIFAR-10 datasets. The training and validation performance is shown to be fairly similar, indicating a low risk of overfitting. We note that the optimal learning rate for the three-dimensional application was found to be , while for the two-dimensional application, it was .
Next, Fig. 3a., and b. present the loss function during training of the explanation optimizer for each class in the 3D and 2D tasks. To train the explanation optimizer a separate hyperparameter tuning, utilizing different learning rates of , , , and , was performed. The light blue shading in Fig. 3a. and b. indicates the min and max loss values, while the solid currve represents the median loss across different learning rates. The best performance for the optimizer was achieved in the three-dimensional application for a learning rate of for the presence of PCS, for the absence of PCS, and for the two-dimensional scenario.
Explainability scores of the explanation optimizer
@@ Results
Fig. 4 presents the faithfulness and complexity scores obtained using various XAI methods alongside our proposed non-linear explanation optimizer for both the two and three-dimensional applications. In terms of faithfulness, our approach consistently outperforms other XAI methods, achieving higher average scores (exceeding 0.20 in 3D and 0.17 in 2D). Regarding complexity, our proposed approach demonstrates lower average values (below 12.15 in 3D and 9.35 in 2D), thus leading to more comprehensible explanations.
@@
Statistical analysis, utilizing ANOVA two-sample tests, revealed highly significant differences between our proposed approach and the alternative XAI methods across all four cases of Fig. 4. Specifically, F values from our analysis were equal to 100.03, 168.02, 18.44, and 147.02 with corresponding F critical values equal to 1.8974, 1.9001, 1.8995, 1.8995 and for 3D faithfulness, 3D complexity, 2D faithfulness, and 2D complexity, respectively. Overall, these results verify that the proposed optimizer yields more optimal explanations than alternative XAI methods both in terms of faithfulness to the deep classifier and in terms of complexity, improving comprehensibility. Thus, our results provide new evidence that optimal explanations are indeed derivable and point to a new direction in XAI applications - that of combining existing methods to achieve higher level understanding of deep networks.
@@
Fig. 5 a.,b.,c. present the explanations in the form of heatmaps, highlighting significant features for both the 3D and 2D applications. In the 3D application, Fig. 5 a., b. present two different examples where the Resnet-18 has successfully identified the absence and existence of PCS in the brain, respectively. The different panels show the important features of the 3D brain (sulci sub-regions in the medial view) that helped the netwotk predict, as identified by different XAI methods and our explanation optimizer. In each case, all important features are presented in the first row, and the top 10% important features are shown in the second row.A significant variability is observed across the explanations provided by existing XAI methods like DeepLift, Kernel Shap, Gradient Shap etc., making it challenging to determine which the optimal explanation is. The explanation from our proposed optimizer (’Explanation Optimizer (non-Linear)’) highlights the sub-regions of the areas where the PCS is usually expected to be located [32, 33] (Fig. 5 a. b). On the contrary, other XAI methods highlight areas that are not as typically related to the PCS connectomics (Gradient Shap; Fig. 5 a.,b. and DeepLift Shap; Fig. 5 b.).
@@
![[Uncaptioned image]](x7.png)
![[Uncaptioned image]](x9.png)
![[Uncaptioned image]](x11.png)
![[Uncaptioned image]](x13.png)
![[Uncaptioned image]](x15.png)
The light blue shading in the panels highlights the min/max and the median loss values obtained during training across various learning rates. (a) The sub-figures present the validation loss of the explanation optimizer for the binary classification task for class 0 (absence of PCS) and class 1 (existence of PCS) in the 3D brain application. (b) The sub-figures present the validation loss of the explanation optimizer for the multi-class classification task for the ten classes in the 2D automobile/animals application.
a
b
c
![[Uncaptioned image]](x18.png)
d
![[Uncaptioned image]](x19.png)
(a) The faithfulness score of different state-of-the-art XAI methods (green color variation) alongside the proposed linear and non-linear explanation optimizer (red color) for the 3D application. (b) Same as (a), but for the complexity score. (c-d) Same as (a-b), but for the 2D application.
a
b
c
![[Uncaptioned image]](x22.png)
d
![[Uncaptioned image]](x23.png)
(a) The panels present the local explanations generated from different state-of-the-art XAI methods, alongside the results from the proposed explanation optimizer (the results of both linear and nonlinear implementations are presented), for the 3D application. Explanations are for a specific human individual, randomly selected from the test cohort of class 0 (absence of PCS). For each XAI method, the top row highlights the feature importance using a color scale from 0 to 1 (blue to red), while the second row highlights the most significant features based on a threshold of 10.0%. (b) Same as (a), but results are for a specific human individual, randomly selected from the test cohort of class 1 (existence of PCS). (c) Same as (a), but for the 2D application. Many different prediction instances are considered (different rows) and only the top 10.0% important features are shown. These explanations are for specific cases randomly selected from the test cohort of the classes: airplane, cat, bird, car, dog, frog, deer. These results are for low resolution (). (d) Same as (c), but the high resolution explanations from our optimizer are presented ().
@@
Thus, the proposed explanation optimizer seems to deliver more realistic patterns than other XAI methods regarding which sub-regions contribute to the identification of the existence or absence of PCS ([32]). Also, the explanation from our optimizer exhibited the highest scores of faithfulness (see also Fig. 4.a), identifying regions of significance that align with results from the study [32]. Last, the proposed approach demonstrated less complexity (see also Fig. 4.b) and focused on specific sub-regions (Fig. 5 a., b.), advancing the understanding of the underlying mechanisms.
@@
For the two-dimensional examples, explanations of many different prediction instances (airplanes, cats, birds, cars, dogs, frogs, and deers) are presented. For simplicity, only the top 10% important features are shown; see Fig. 5 c. The results suggest that our proposed method provides more accurate explanations of the network’s predictions compared to other XAI methods, emphasizing objects and foreground parts that are expected to be critical for classification. For instance, in the first airplane example (top row), saliency and DeepLift primarily focus on the airplane’s head, Kernel Shap focuses on the tail, and Gradient Shap focuses on the body and head, while our proposed optimizer identifies all parts of the air plane as significant features. Similar remarks are valid for the instances of the car classification, where the optimal explanation very explicitly identifies all the body parts of the car as important features to the network prediction.
Why utilizing a non-linear optimizer structure over linear alternatives?
@@
A question that may arise is why we opted to use a non-linear structure in our proposed framework (see Fig. 2 and Fig. 6) to reconstruct the unique and optimal explanation for the deep learning network, instead of using a linear structure. To investigate this further, we explored a linearly structured optimizer (see Supplementary Fig. 1) and compared the results with the non-linear one. Specifically, in the 3D neuroscience application, where the complexity of the space is higher, we decided to further examine reconstruction structures beyond the non-linear approach, namely, encoder, bottleneck, and decoder with non-linear activations and blocks (non-linear structure, Fig. 6). Therefore, we employed a linear structure with attention mechanisms to estimate scalable weights to be used in a weighted average of the existing XAI methodologies (referred to in the subsection ’Deep learning architecture’ of the Methods section). The baseline ’weighted average’ mentioned in the Methods section and presented in Fig. 6 is considered as another linear approach employing fixed weights. Both linear approaches, whether employing fixed weights (’Weighted Average’) or trainable weights (’Explanation Optimizer (Linear)’), yielded comparable results to the existing XAI methods, as indicated in Fig. 4 a. and b. This suggests that linear solutions may not significantly outperform existing approaches in optimizing explanations for deep learning networks. Consequently, there is a rationale for exploring non-linear approaches, as done in this study. The results of a sensitivity analysis comparing the linear and non-linear structures for different learning rates are provided in the Supplementary material section, entitled ’Sensitivity analysis for linear and non-linear structures’ (see Supplementary Fig. 2). The main observation is that the linear structure is more constrained in faithfulness and complexity scores during training, ultimately converging to lower faithfulness and higher complexity scores compared to the non-linear structure. In contrast, the nonlinear structure exhibits a more stochastic behaviour during training, ultimately converging to a significantly more optimal solution.
Justifying the pursuit of optimal explanations for deep learning networks
@@
In this study, we propose a method to address the challenge of identifying the optimal explanation among various XAI methods. We acknowledge in the manuscript that while adopting a holistic approach and using multiple XAI methods has been shown to be beneficial in some settings ([14, 17, 16]), it can result in highly different explanations, sometimes yielding overwhelming and incomprehensible outcomes, thus, lowering the trust to the network’s predictions. Our study demonstrates that opting to derive optimal explanations based on user-defined explainability quality metrics is feasible (see Fig. 4) and a promising alternative. Visual examples of explanations validate this claim, reinforcing the credibility of such an approach (see Fig. 5).
@@
Deriving optimal explanations has been successfully achieved here by utilizing an innovative loss function. As a first component of the loss, we used the two scores of faithfulness and complexity, so as to ensure that the derived explanation is faithful to the network’s prediction-making process and maximally comprehensible. Additionally, we used a third term in the loss representing the similarity between the derived explanation and a baseline weighted average of explanations from existing XAI methods. By employing this similarity term, we further constrained the training of the optimizer, facilitated convergence and prevented divergence towards outlier and/or trivial solutions (e.g., red noise, impulse noise, etc.). By adjusting the importance of each one of these loss terms, one may formulate a unique cost function tailored to solve the problem of optimal explanation based on the needs of their application. We note that in our study, we chose the relative importance of these loss terms via hyper-parameter tuning (see Methods).
@@
The final component of the proposed framework involved an Up-Sampling layer, aiming to provide a higher resolution explanation than the resolution of the prediction task (Fig. 6). This higher resolution explanation is vital, particularly in the domain of neuroscience, as it allows for the capture of more intricate details, enhancing the accuracy and comprehensiveness of the explanation. Fig. 5a. and b. show that the results from the ’Explanation Optimizer (non-Linear) high-resolution’ are slightly different from compared to the results from the ’Explanation Optimizer (non-Linear)’. The high resolution explanations are more in alignment with the given protocol for the identification of PCS ([33, 32]), supporting that the Up-sampling layer is a beneficial addition to our approach.
Beyond the two and three dimensions
@@
We demonstrate our proposed framework in both 3D and 2D applications (Fig. 2a. and b.). For the sake of generalization of our proposed framework, we extended it to ’M’ dimensions (i.e., referring to the input space), based on the principles outlined in our approach (see Fig. 6). This generalized version can be adapted to any Artificial Intelligence application. Specifically, we denote ’K’ different XAI methods used as a baseline, with the ’input token’ representing the input features in a specific AI task of interest (see Fig. 6).
The generalized version of the proposed explanation optimizer in M dimensions.M: the number of dimensionality of the input space, K: the number of XAI methods used as baseline, input token: the input features for the AI task, XAI token: The local explanations from the baseline methods, : The assigned weights based on the scoring of each XAI method on different explainability metrics (here complexity and faithfulness), adjusted token: The weighted product of and the XAI token, Weight Explanation Token (WET): the weighted average based on the K adjusted tokens of XAI methods, the High Resolution Token (HRT) represents the high resolution version of the optimal explanation after the ’non-linear structure’ layer. Low Resolution Token (LRT) represents the low resolution that is the resolution of the input token.
@@
The ’XAI token’ represents the essential features needed for the local explanation, generated by the baseline XAI methods. Furthermore, we denote different assigned weights derived from explainability metrics such as complexity and faithfulness, and we use the term ’adjusted token’ to refer to the products of the weights with the XAI tokens. We define the ’Weighted Explanation Token (WET)’ as the weighted-average product derived from the adjusted tokens of the XAI methods. Additionally, the ’High Resolution Token (HRT)’ represents the high-resolution space post the ’non-linear structure’ layer in the explanation optimization process, while the ’Low Resolution Token (LRT)’ denotes the input token’s low-resolution space.
@@ Discussion
Our study explores the degree to which such an explanation is even derivable. Our proposed framework, inspired by Herbert Simon’s seminal work on ”bounded rationality” and ”satisficing” decision-making, is shown to be able to derive such a unique and optimal explanation for pre-trained deep learning networks, tailored to specific computer vision tasks. By integrating diverse explanations from established XAI methods, employing a non-linear network architecture and an innovative loss function, we achieve to reconstruct a unique explanation that faithfully captures the essence of the network’s decision-making process and minimizes complexity. To validate our framework, we conducted experiments on multi-class and binary classification tasks in both two- and three-dimensional spaces, focusing on automobiles-animals and neuroscience imaging domains.
Our results demonstrate the effectiveness of the framework in achieving high faithfulness and low complexity in explanations. Statistical analysis revealed significant differences between our approach and alternative state-of-the-art methods, further affirming its efficacy. Visual explanations provided by our framework offer insights into the decision-making process of deep learning models. In two-dimensional scenarios, our approach consistently outperformed other XAI methods, emphasizing significant features crucial for classification. In three-dimensional applications, our framework showcased less complexity and focused on specific areas, advancing our understanding of underlying mechanisms. Our study also discusses the rationale behind favoring non-linear optimizing structures over linear alternatives. While linear solutions demonstrate comparable performance to the state-of-the-art XAI methods, non-linear approaches offer greater flexibility and accuracy in optimizing a highly faithful and less complex explanation for deep learning networks. Our optimizer is successful in part because of the utilization of a unique cost function that consists of explanability scores and a similarity score that constraints the convergence of the optimizer and avoids reaching non meaningful, trivial solutions during training. The generalized version of the proposed framework can be used in any explanation task of a N-dimensional space.

### v2

@@ 4.4 Results of classification training and optimization of explanations
4.5 Explainability scores of the explanation optimizer
@@
Fig. 5 and Supplementary material Fig. 5 present the faithfulness and complexity scores obtained using various XAI methods alongside our proposed non-linear explanation optimizer for the three-dimensional and two-dimensional applications, respectively. In terms of faithfulness, our approach consistently outperforms other XAI methods, achieving higher average scores (exceeding 0.20 in the 3D application and 0.17 in the 2D application). Regarding complexity, our proposed approach demonstrates lower average values (below 12.15 in the 3D application and 9.34 in the 2D application), thus leading to more comprehensible explanations. Regarding the 2D application, we briefly mention that the most faithful XAI method following our approach is DeepLiftShap, achieving an average score of 0.07, while the simplest XAI method behind our optimizer is KernelShap, with an average complexity score of 9.38.
@@
We have performed statistical analysis (see Figure 11 and section 5 in Supplementary material) using a two-tailed ANOVA and normality tests across four different scenarios. The ANOVA results show highly significant differences between groups, with F-values of 147, 17.01, 1142, and 37.35, and corresponding p-values all less than 2e-16, indicating that the group differences are statistically significant. However, the Shapiro-Wilk tests indicate significant deviations from normality, with p-values below 2.2e-16 in some cases, suggesting that the normality assumption for ANOVA may not hold. To address this concern, a nonparametric Kruskal-Wallis test was also conducted, yielding p-values of 3.3e-65, 3.22e-18, 9.32e-22, and 3.29e-18, which also support the conclusion of significant differences between groups, despite the non-normality in the data. Based on these results, we can confirm the significant improved performance of our proposed explanation optimizer across all the different experiments. Overall, these results verify that the proposed optimizer yields more optimal explanations than alternative XAI methods both in terms of faithfulness to the deep classifier and in terms of complexity, improving comprehensibility. Thus, our results provide new evidence that optimal explanations are indeed derivable and point to a new direction in XAI applications - that of combining existing methods to achieve higher level understanding of deep networks.
@@
Fig. 6 and Supplementary material Fig. 6 a, present specific explanations in the form of heatmaps, highlighting significant features for both the 3D and 2D applications. In the 3D application, Fig. 6 a., b. present two different examples where the Resnet-18 has successfully identified the absence and existence of PCS in the brain, respectively. The different panels show the important features of the 3D brain (sulci sub-regions in the medial view) that helped the netwotk predict, as identified by different XAI methods and our explanation optimizer. In each case, all important features are presented in the first row, and the top 10% important features are shown in the second row. A significant variability is observed across the explanations provided by existing XAI methods like DeepLift, Kernel Shap, Gradient Shap etc., making it challenging to determine which the optimal explanation is. The explanation from our proposed optimizer (’Explanation Optimizer (non-Linear)’) highlights the sub-regions of the areas where the PCS is usually expected to be located (Mamalakis et al., 2024b; Mitchell et al., 2023) (Fig. 6 a. b). On the contrary, other XAI methods highlight areas that are not as typically related to the PCS connectomics (Gradient Shap; Fig. 6 a.,b. and DeepLift Shap; Fig. 6 b.).
@@
Thus, the proposed explanation optimizer seems to deliver more realistic patterns than other XAI methods regarding which sub-regions contribute to predicting the existence or absence of PCS (Mamalakis et al., 2024b). Also, the explanation from our optimizer exhibited the highest scores of faithfulness (see also Fig. 5.a), identifying regions of significance that align with results from the study (Mamalakis et al., 2024b). Last, the proposed approach demonstrated less complexity (see also Fig. 5.b) and focused on specific sub-regions (Fig. 6 a., b.), advancing the understanding of the underlying mechanisms.
@@
For the two-dimensional examples, explanations of many different prediction instances (airplanes, cats, birds, cars, dogs, frogs, and deers) are presented. For simplicity, only the top 10% important features are shown; see Supplementary material Fig. 6 a. The results suggest that our proposed method provides more accurate explanations of the network’s predictions compared to other XAI methods, emphasizing objects and foreground parts that are expected to be critical for classification. For instance, in the first airplane example (top row), saliency and DeepLift primarily focus on the airplane’s head, Kernel Shap focuses on the tail, and Gradient Shap focuses on the body and head, while our proposed optimizer identifies all parts of the air plane as significant features. Similar remarks are valid for the car classification, where the optimal explanation very explicitly identifies all the body parts of the car as important features to the network prediction.
4.6 Why utilizing a non-linear optimizer structure over linear alternatives?
@@
A question that may arise is why we opted to use a non-linear structure in our proposed framework (see Fig. 2 and Fig. 7) to reconstruct the unique and optimal explanation for the deep learning network, instead of using a linear structure. To investigate this further, we explored a linearly structured optimizer (see Supplementary material Fig. 2) and compared the results to those of the non-linear one. Specifically, in the 3D neuroscience application, where the complexity of the space is higher, we decided to further examine reconstruction structures beyond the non-linear approach. Therefore, we employed a linear structure with attention mechanisms to estimate scalable weights to be used in a weighted average of the existing XAI methodologies. The baseline ’weighted average’ mentioned in the Methods section and presented in Fig. 7 is considered as another linear approach employing fixed weights. Both linear approaches, whether employing fixed weights (’Weighted Average’) or trainable weights (’Explanation Optimizer (Linear)’), yielded comparable results to the existing XAI methods, as indicated in Fig. 5 a. and b. This suggests that linear solutions may not significantly outperform existing approaches in optimizing explanations for deep learning networks. Consequently, there is a rationale for exploring non-linear approaches, as done in this study.
A sensitivity analysis comparing the linear and non-linear structures for different learning rates is provided in the Supplementary material, entitled ’Sensitivity analysis for linear and non-linear structures’ (see Supplementary Fig. 8). The main result is that the linear structure is more constrained in faithfulness and complexity scores during training, ultimately converging to lower faithfulness and higher complexity scores compared to the non-linear structure. In contrast, the nonlinear structure exhibits a more stochastic behaviour during training, ultimately converging to a significantly more optimal solution.
4.7 Justifying the pursuit of optimal explanations for deep learning networks
@@
In this study, we propose a method to address the challenge of identifying the optimal explanation among various XAI methods. We acknowledge in the manuscript that while adopting a holistic approach and using multiple XAI methods has been shown to be beneficial in some settings (Mamalakis et al., 2022; Dikshit et al., 2024; McGovern et al., 2019), it can result in highly different explanations, sometimes yielding overwhelming and incomprehensible outcomes, thus, lowering the trust to the network’s predictions. Our study demonstrates that opting to derive optimal explanations based on user-defined explainability quality metrics is feasible (see Fig. 5 and Supplementary material Fig. 5) and a promising alternative. Visual examples of explanations validate this claim, reinforcing the credibility of such an approach (see Fig. 6 and Supplementary material Fig. 6).
@@
Deriving optimal explanations has been successfully achieved here by utilizing an innovative loss function. As a first component of the loss, we used the two scores of faithfulness and complexity, so as to ensure that the derived explanation is faithful to the network’s prediction-making process and maximally comprehensible. Additionally, we used a third term in the loss representing the similarity between the derived explanation and a baseline weighted average of explanations from existing XAI methods. By employing this similarity term, we further constrained the optimizer’s training, promoted convergence, and mitigated divergence toward outlier or trivial solutions (e.g., red noise, impulse noise). By adjusting the importance of each one of these loss terms, one may formulate a unique cost function tailored to solve the problem of optimal explanation based on the needs of their application. We note that in our study, we chose the relative importance of these loss terms via hyper-parameter tuning (see Methods).
@@
The final component of the proposed framework involved an Up-Sampling layer, aiming to provide a higher resolution explanation than the resolution of the prediction task (Fig. 7). This higher resolution explanation is vital, particularly in the domain of neuroscience, as it allows for the capture of more intricate details, enhancing the accuracy and comprehensiveness of the explanation. Fig. 6a. and b. show that the results of the ’Explanation Optimizer (non-Linear) high-resolution’ are slightly different from the results of the ’Explanation Optimizer (non-Linear)’. The high resolution explanations are more in alignment with the given protocol for the identification of PCS (Mitchell et al. (2023); Mamalakis et al. (2024b)), supporting that the Up-sampling layer is a beneficial addition to our approach.
4.8 Beyond the two and three dimensions
@@
We demonstrate our proposed framework in both 3D and 2D applications (Fig. 2 and Supplementary material Fig. 1, respectively). For the sake of generalization of our proposed framework, we extended it to ’M’ dimensions (i.e., referring to the input space), based on the principles outlined in our approach (see Fig. 7). This generalized version can be adapted to any Artificial Intelligence application. Specifically, we denote ’K’ different XAI methods used as a baseline, with the ’input token’ representing the input features in a specific AI task of interest (see Fig. 7).
@@
The ’XAI token’ represents the essential features needed for the local explanation, generated by the baseline XAI methods. Furthermore, we denote different assigned weights derived from explainability metrics such as complexity and faithfulness, and we use the term ’adjusted token’ to refer to the products of the weights with the XAI tokens. We define the ’Weighted Explanation Token (WET)’ as the weighted-average derived from the adjusted tokens of the XAI methods. Additionally, the ’High Resolution Token (HRT)’ represents the high-resolution space post the ’non-linear structure’ layer in the explanation optimization process, while the ’Low Resolution Token (LRT)’ denotes the input token’s low-resolution space.
@@ 5 Discussion
Our study explores the degree to which such an explanation is even derivable. Our proposed framework, inspired by Herbert Simon’s seminal work on "bounded rationality" and "satisficing" decision-making, is shown to be able to derive such a unique and optimal explanation for pre-trained deep learning networks, tailored to specific computer vision tasks. By integrating diverse explanations from established XAI methods, employing a non-linear network architecture and an innovative loss function, we achieve to reconstruct a unique explanation that faithfully captures the essence of the network’s decision-making process and minimizes complexity. To validate our framework, we conducted experiments on multi-class and binary classification tasks in both two- and three-dimensional spaces, focusing on automobiles-animals and neuroscience imaging domains.
Our results demonstrate the effectiveness of the framework in achieving high faithfulness and low complexity in explanations. Statistical analysis revealed significant differences between our approach and alternative state-of-the-art methods, further affirming its efficacy. Visual explanations provided by our framework offer insights into the decision-making process of deep learning models. In two-dimensional scenarios, our approach consistently outperformed other XAI methods, emphasizing significant features crucial for classification. In three-dimensional applications, our framework showcased less complexity and focused on specific areas, advancing our understanding of underlying mechanisms. Our study also discusses the rationale behind favoring non-linear optimizing structures over linear alternatives. Importantly, our results highlight that a straightforward linear combination of existing XAI methods does not improve faithfulness or reduce complexity meaningfully, as demonstrated in Figure 5 (’Weighted Average’ and ’Explanation Optimizer (linear)’). In contrast, our non-linear explanation optimizer achieves significant advancements, reinforcing the non-trivial nature of designing such a unifying framework. Our optimizer is successful in part because of the utilization of a unique cost function that consists of explanability scores and a similarity score that constraints the convergence of the optimizer and avoids reaching non meaningful, trivial solutions during training. The generalized version of the proposed framework can be used in any explanation task of a multi-dimensional space.